# Quantum Curriculum Learning

### v2

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to achieve quantum advantage. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. We define the curriculum criteria based on the data density ratio between tasks to determine the curriculum order. We also implement a dynamic learning schedule to emphasize the significance of quantum data in optimizing the loss function. Empirical evidence shows that Q-CurL significantly enhances the training convergence and the generalization for unitary learning tasks and improves the robustness of quantum phase recognition tasks. Our framework provides a general learning strategy, bringing QML closer to realizing practical advantages.
@@
Introduction.— In the emerging field of quantum computing (QC), there is promise in using quantum computers at sufficiently large scales to solve certain machine learning (ML) problems much more efficiently than classical methods. This synergy between ML and QC has led to the creation of a new field known as quantum machine learning (QML) [1, 2], even though the practicality of its applications remains unclear. There is a question as to whether speed is the only metric by which QML algorithms should be judged [3]. In fact, classical ML focuses on identifying and reproducing features from data statistics. The initial hope is that QML can detect correlations in the classical data or generate new patterns that would be difficult for classical algorithms to achieve [4, 5, 6, 7, 8]. However, there is no clear evidence that analyzing classical data inherently requires quantum effects. This suggests a fundamental shift in the research community’s perspective: it is preferable to use QML on data that is already quantum in nature [9, 10, 11, 12, 13, 14].
@@
The typical learning process in QML involves extensive exploration within the domain landscape of a loss function. This function measures the discrepancy between the quantum model’s predictions and the actual values, aiming to locate its minimum using classical computers. However, the optimization often encounters pitfalls such as getting trapped in local minima [15, 16] or barren plateau regions [17]. These scenarios require substantial quantum computational resources to navigate the loss landscape successfully. Additionally, improving task accuracies necessitates evaluating numerous model configurations, especially against extensive datasets. Given the limitation of current quantum resources, this poses a considerable challenge. Therefore, in designing QML models, we must focus not only on their architectural aspects but also on efficient learning strategies.
@@
The perspective of quantum resources refocuses our attention on the concept of learning. In ML, learning refers to the process through which a computer system enhances its performance on a specific task over time by acquiring and integrating knowledge or patterns from data. We can improve current QML algorithms by making this process more efficient. For example, curriculum learning [18] is inspired by the human learning process, based on the intuitive observation that we often begin with simpler concepts before progressing to more complex ones. This insight leads to the development of a strategy for sampling or task scheduling—a curriculum—that introduces simpler samples or tasks to the model before moving on to more challenging ones. Although curriculum learning has been extensively applied and studied in classical ML [19, 20, 21], its exploration in the QML field, especially regarding quantum data, is still in the early stages. The most relevant research has focused on investigating model transfer learning within hybrid classical-quantum neural networks [22]. Typically, this involves starting with a pre-trained classical network that is then modified and enhanced by adding a variational quantum circuit at its final layer. Despite the potential benefits, there is still a lack of concrete evidence that effectively using a curriculum learning framework to schedule tasks and samples improves QML techniques.
In this Letter, we explore the potential of curriculum learning within the context of QML using quantum data. We demonstrate the feasibility of implementing curriculum learning in a quantum framework called quantum curriculum learning (Q-CurL). We consider two common scenarios in training QML models. First, a main task, which may be challenging or limited by data availability, can be facilitated through the preparatory parameter adjustment of an auxiliary task. This auxiliary task is comparatively easier or more data-rich. However, it is necessary to establish the criteria that make an auxiliary task beneficial for a main task. Second, quantum data may carry different weights of importance in the optimization process. Focusing on the partial loss function of difficult data could lead to poorer generalization. We propose two principal approaches for these scenarios: task-based Q-CurL [Fig.1(a)], which determines the curriculum order based on the data density ratio between tasks, and data-based Q-CurL [Fig.1(b)], which employs a dynamic learning schedule that adjusts data weights, thereby prioritizing the importance of quantum data in the optimization. Empirical evidence demonstrates the utility of Q-CurL in enhancing training convergence and generalization when learning unitary dynamics. Furthermore, we show that data-based Q-CurL contributes to increased robustness in learning. Specifically, in scenarios with noisy labels, it can prevent complete memorization of the training data, thereby avoiding overfitting and enhancing generalization performance.
@@
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis, prediction model) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$ such that $h({\bm{x}})\approx f({\bm{x}})$. To evaluate the correctness of the hypothesis $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find a hypothesis $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to model $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data [23].
@@
We consider the unitary learning task to demonstrate the curriculum criteria based on $c_{M,m}$. We aim to optimize the parameters ${\bm{\theta}}$ of a $Q$-qubit parameterized quantum circuit $U({\bm{\theta}})$, such that, for the optimized parameters ${\bm{\theta}}_{\textrm{opt}}$, $U({\bm{\theta}}_{\textrm{opt}})$ can approximate an unknown $Q$-qubit unitary $V$ ($U,V\in\mathcal{U}(\mathbb{C}^{2^{Q}})$).
@@
The parameterized ansatz unitary $U({\bm{\theta}})$ can be modeled as $U({\bm{\theta}})=\prod_{l=1}^{L}U^{(l)}({\bm{\theta}}_{l})$, consisting of $L$ repeating layers of unitaries. Each layer $U^{(l)}({\bm{\theta}}_{l})=\prod_{k=1}^{K}\exp{(-i\theta_{lk}H_{k})}$ is composed of $K$ unitaries, where $H_{k}$ are Hermitian operators, ${\bm{\theta}}_{l}$ is a $K$-dimensional vector, and ${\bm{\theta}}=\{{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{L}\}$ is the $LK$-dimensional parameter vector.
@@
Data-based Q-CurL.— We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Apart from improving generalization, the benefit of data-based Q-CurL lies in its resistance to noise, which is especially needed in QML. Existing QML models can accurately fit partially corrupted labels to quantum states in the training data but fail on the test data [25]. We show that data-based Q-CurL can enhance the robustness based on the dynamic weighting of the difficulty fitting quantum data to corrupted labels.
@@
Here, ${\bm{w}}=(w_{1},\ldots,w_{N})$, $\ell_{i}=\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})$, and $w^{2}_{i}$ is the regularization term controlled by the hyper-parameter $\gamma>0$. The threshold $\eta$ distinguishes easy and hard samples with $e^{w_{i}}$ emphasizing the loss $l_{i}\ll\eta$ (easy sample) and neglecting the loss $l_{i}\gg\eta$ (hard samples, such as data with corrupted labels). The minimization problem is reduced to $\textup{min}_{{\bm{\theta}}}\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$, where ${\bm{\theta}}$ is the parameter of the hypothesis $h$. Here, $\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$ is decomposed at each loss $\ell_{i}$ and solved without quantum resources as $w_{i}=\textup{argmin}_{w}(l_{i}-\eta)e^{w}+\gamma w^{2}$. To control the difficulty of the samples, in each training epoch, we set $\eta$ as the average value of all $\ell_{i}$ obtained from the previous epoch. Therefore, $\eta$ changes dynamically during the early stages of training but remains almost constant during the convergence periods.
@@
We apply the data-based Q-CurL to the quantum phase recognition task investigated in Ref. [10] to demonstrate that it can improve the generalization of the learning model. Here, we consider a one-dimensional cluster Ising model with open boundary conditions, whose Hamiltonian with $Q$ qubits is given by $H=-\sum_{j=1}^{Q-2}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}-h_{1}\sum_{j=1}^{Q}\sigma^{x}_{j}-h_{2}\sum_{j=1}^{Q-1}\sigma^{x}_{j}\sigma^{x}_{j+1}.$ Depending on the coupling constants $(h_{1},h_{2})$, the ground state wave function of this Hamiltonian can exhibit multiple states of matter, such as the symmetry-protected topological phase, the paramagnetic state, and the anti-ferromagnetic state. We employ the quantum convolutional neural network (QCNN) model [10] with binary cross-entropy loss for training. Without Q-CurL, we use the conventional loss $\hat{R}(h)=(1/N)\sum_{i=1}^{N}\ell_{i}$ for the training and test phase. In data-based Q-CurL, we train the QCNN with the loss $\hat{R}(h,{\bm{w}})$ while using $\hat{R}(h)$ to evaluate the generalization on the test data set. We use 40 and 400 ground state wave functions for the training and test phase, respectively (see [23] for details of settings).
@@
To evaluate the effectiveness of the data-based Q-CurL in considering the difficulty of the data in training, we consider the scenario of fitting corrupted labels. Given a probability $p$ ($0\leq p\leq 1$) representing the noise level, the true label $y_{i}\in\{0,1\}$ of training state |$\ket{\psi_{i}}$ is transformed to the label $1-y_{i}$ with probability $p$, while it remains the true label with probability $1-p$. Figure 3 illustrates the performance of trained QCNN on test data across different noise levels. There is no significant difference at low noise levels, but as the noise level increases, the conventional training fails to generalize effectively. In this case, introducing data-based Q-CurL in training (red lines) reduces the test loss and enhances testing accuracy compared to the conventional method (blue lines).

### v3

@@ Abstract
Quantum machine learning (QML) requires significant quantum resources to address practical real-world problems. When the underlying quantum information exhibits hierarchical structures in the data, limitations persist in training complexity and generalization. Research should prioritize both the efficient design of quantum architectures and the development of learning strategies to optimize resource usage. We propose a framework called quantum curriculum learning (Q-CurL) for quantum data, where the curriculum introduces simpler tasks or data to the learning model before progressing to more challenging ones. Q-CurL exhibits robustness to noise and data limitations, which is particularly relevant for current and near-term noisy intermediate-scale quantum devices. We achieve this through a curriculum design based on quantum data density ratios and a dynamic learning schedule that prioritizes the most informative quantum data. Empirical evidence shows that Q-CurL significantly enhances training convergence and generalization for unitary learning and improves the robustness of quantum phase recognition tasks. Q-CurL is effective with broad physical learning applications in condensed matter physics and quantum chemistry.
@@
Introduction.— In the emerging field of quantum computing (QC), there is potential to use large-scale quantum computers to solve certain machine learning (ML) problems far more efficiently than classical methods. This synergy between ML and QC has given rise to quantum machine learning (QML) [1, 2], although its practical applications remain uncertain. Classical ML traditionally focuses on extracting and replicating features based on data statistics, while QML is hoped to detect correlations in classical data or generate patterns that are challenging for classical algorithms to achieve [3, 4, 5, 6, 7]. However, it remains unclear whether analyzing classical data fundamentally requires quantum effects. Furthermore, there is a question as to whether speed is the only metric by which QML algorithms should be judged [8]. This suggests a fundamental shift: it is preferable to use QML on data that is already quantum in nature [9, 10, 11, 12, 13, 14].
@@
The learning process in QML involves extensive exploration within the domain landscape of a loss function. This function measures the discrepancy between the quantum model’s predictions and the actual values, aiming to locate its minimum. However, the optimization often encounters pitfalls such as getting trapped in local minima [15, 16] or barren plateau regions [17]. These scenarios require substantial quantum resources to navigate the loss landscape successfully. Additionally, improving accuracies necessitates evaluating numerous model configurations, especially against extensive datasets. Given the limitation of quantum resources in designing QML models, we must focus not only on their architectural aspects but also on efficient learning strategies.
@@
The perspective of quantum resources refocuses our attention on the concept of learning. In ML, learning refers to the process through which a computer system enhances its performance on a specific task over time by acquiring and integrating knowledge or patterns from data. We can improve current QML algorithms by making this process more efficient. For example, curriculum learning [18], inspired by human learning, builds on the idea of introducing simpler concepts before progressing to complex ones, forming a strategy—a curriculum—that presents easier samples or tasks first. Although curriculum learning has been extensively applied in classical ML [19, 20, 21], its exploration in the QML field, especially regarding quantum data, is still in the early stages. Existing research has primarily examined model transfer learning in hybrid classical-quantum networks [22], where a pre-trained classical model is enhanced by adding a variational quantum circuit. However, there is still limited evidence showing that curriculum learning can effectively improve QML by scheduling tasks and samples.
We explore the potential of curriculum learning using quantum data. We implement a quantum curriculum learning (Q-CurL) framework in two common scenarios. First, a main quantum task, which may be challenging due to the high-dimensional nature of the parameter space or the limitation of data availability, can be facilitated through the hierarchical parameter adjustment of auxiliary tasks. These auxiliary tasks are comparatively easier or more data-rich. However, it is necessary to establish the criteria that make an auxiliary task beneficial for a main task. Second, QML often involves noisy inputs that exhibit a hierarchical arrangement of entanglement or noisy labels, reflecting levels of importance during the optimization process. Recognizing these levels is essential for ensuring the robustness and reliability of QML methods in practical scenarios.
We propose two principal approaches to address the outlined scenarios: task-based Q-CurL [Fig. 1(a)] for the first and data-based Q-CurL [Fig. 1(b)] for the second scenario. In task-based Q-CurL, the curriculum order is defined by the fidelity-based kernel density ratio between quantum datasets. This enables efficient auxiliary task selection without solving each one, reducing data demands for the main task and decreasing training epochs, even if total data requirements stay constant. In data-based Q-CurL, we employ a dynamic learning schedule that adjusts data weights to prioritize quantum data in optimization. This adaptive cost function is broadly applicable to any cost function without requiring additional quantum resources. Empirical evidence shows that task-based Q-CurL enhances training convergence and generalization when learning complex unitary dynamics. Additionally, data-based Q-CurL increases robustness, particularly in noisy-label scenarios, by preventing complete memorization of the training data. This avoids overfitting and improves generalization in the quantum phase detection task. These results suggest that Q-CurL could be broadly effective for physical learning applications.
@@
Task-based Q-CurL.— We formulate a framework for task-based Q-CurL. The target of learning is to find a function (or hypothesis) $h:{\mathcal{X}}\to{\mathcal{Y}}$ within a hypothesis set ${\mathcal{H}}$ that approximates the true function $f$ mapping ${\bm{x}}\in{\mathcal{X}}$ to ${\bm{y}}=f({\bm{x}})\in{\mathcal{Y}}$. To evaluate the correctness of $h$ given the data $({\bm{x}},{\bm{y}})$, the loss function $\ell:{\mathcal{Y}}\times{\mathcal{Y}}\to\mathbb{R}$ is used to measure the approximation error $\ell(h({\bm{x}}),{\bm{y}})$ between the prediction $h({\bm{x}})$ and the target ${\bm{y}}$. We aim to find $h\in{\mathcal{H}}$ that minimizes the expected risk over the distribution $P({\mathcal{X}},{\mathcal{Y}})$:
@@
The curriculum weight $c_{M,m}$ can be determined using the density ratio $r({\bm{x}},{\bm{y}})=\dfrac{p^{(M)}({\bm{x}},{\bm{y}})}{p^{(m)}({\bm{x}},{\bm{y}})}$ without requiring the density estimation of $p^{(M)}({\bm{x}},{\bm{y}})$ and $p^{(m)}({\bm{x}},{\bm{y}})$. The key idea is to estimate $r({\bm{x}},{\bm{y}})$ using a linear model $\hat{r}({\bm{x}},{\bm{y}}):={\bm{\alpha}}^{\top}{\bm{\phi}}({\bm{x}},{\bm{y}})=\sum_{i=1}^{N_{M}}\alpha_{i}\phi_{i}({\bm{x}},{\bm{y}}),$ where the vector of basis functions is ${\bm{\phi}}({\bm{x}},{\bm{y}})=(\phi_{1}({\bm{x}},{\bm{y}}),\ldots,\phi_{N_{M}}({\bm{x}},{\bm{y}}))$, and the parameter vector ${\bm{\alpha}}=(\alpha_{1},\ldots,\alpha_{N_{M}})^{\top}$ is learned from data [23].
@@
We consider the unitary learning task to verify the curriculum criteria based on $c_{M,m}$. We aim to optimize the parameters ${\bm{\theta}}$ of a $Q$-qubit circuit $U({\bm{\theta}})$, such that, for the optimized parameters ${\bm{\theta}}_{\textrm{opt}}$, $U({\bm{\theta}}_{\textrm{opt}})$ can approximate an unknown $Q$-qubit unitary $V$ ($U,V\in\mathcal{U}(\mathbb{C}^{2^{Q}})$).
@@
The parameterized ansatz $U({\bm{\theta}})$ can be modeled as $U({\bm{\theta}})=\prod_{l=1}^{L}U^{(l)}({\bm{\theta}}_{l})$, consisting of $L$ repeating layers of unitaries. Each layer $U^{(l)}({\bm{\theta}}_{l})=\prod_{k=1}^{K}\exp{(-i\theta_{lk}H_{k})}$ is composed of $K$ unitaries, where $H_{k}$ are Hermitian operators, ${\bm{\theta}}_{l}$ is a $K$-dimensional vector, and ${\bm{\theta}}=\{{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{L}\}$ is the $LK$-dimensional parameter vector.
@@
Data-based Q-CurL.— We present a form of data-based Q-CurL that dynamically predicts the easiness of each sample at each training epoch, such that easy samples are emphasized with large weights during the early stages of training and conversely. Remarkably, it does not involve pre-training or additional training data, thereby avoiding any increase in quantum resource requirements.
Apart from improving generalization, data-based Q-CurL offers resistance to noise. This feature is particularly valuable in QML, where clean annotated data are often costly while noisy data are abundant. Existing QML models can accurately fit corrupted labels in the training data but often fail on test data [25]. We demonstrate that data-based Q-CurL enhances robustness by dynamically weighting the difficulty of fitting corrupted labels.
@@
Here, ${\bm{w}}=(w_{1},\ldots,w_{N})$, $\ell_{i}=\ell(h({\bm{x}}_{i}),{\bm{y}}_{i})$, and $w^{2}_{i}$ is the regularization term controlled by the hyper-parameter $\gamma>0$. The threshold $\eta$ distinguishes easy and hard samples with $e^{w_{i}}$ emphasizing the loss $l_{i}\ll\eta$ (easy sample) and neglecting the loss $l_{i}\gg\eta$ (hard samples, such as data with corrupted labels) [^1] In Supplementary Material, we have also discussed an interesting scenario where the modified loss in Eq. (7) can be used to emphasize complex quantum data during training, potentially reducing generation errors in quantum phase detection tasks under specific conditions. This aligns with the numerical results reported in Ref. [26], which appeared on arXiv after our paper.. The optimization is reduced to $\textup{min}_{{\bm{\theta}}}\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$, where ${\bm{\theta}}$ is the parameter of the hypothesis $h$. Here, $\textup{min}_{{\bm{w}}}\hat{R}(h,{\bm{w}})$ is decomposed at each loss $\ell_{i}$ and solved without quantum resources as $w_{i}=\textup{argmin}_{w}(l_{i}-\eta)e^{w}+\gamma w^{2}$. To control the difficulty of the samples, in each training epoch, we set $\eta$ as the average value of all $\ell_{i}$ obtained from the previous epoch. Therefore, $\eta$ adjusts dynamically in the early training stages but stabilizes near convergence.
@@
We apply the data-based Q-CurL to the quantum phase recognition task investigated in Ref. [10] to demonstrate that it can improve the generalization of the learning model. Here, we consider a one-dimensional cluster Ising model with open boundary conditions, whose Hamiltonian with $Q$ qubits is given by $H=-\sum_{j=1}^{Q-2}\sigma^{z}_{j}\sigma^{x}_{j+1}\sigma^{z}_{j+2}-h_{1}\sum_{j=1}^{Q}\sigma^{x}_{j}-h_{2}\sum_{j=1}^{Q-1}\sigma^{x}_{j}\sigma^{x}_{j+1}.$ Depending on the coupling constants $(h_{1},h_{2})$, the ground state wave function of this Hamiltonian can exhibit multiple states of matter, such as the symmetry-protected topological phase, the paramagnetic state, and the anti-ferromagnetic state. We employ the quantum convolutional neural network (QCNN) model [10] with binary cross-entropy loss for training. Without Q-CurL, we use the conventional loss $\hat{R}(h)=(1/N)\sum_{i=1}^{N}\ell_{i}$ for the training and test phase. In data-based Q-CurL, we train the QCNN with the loss $\hat{R}(h,{\bm{w}})$ while using $\hat{R}(h)$ to evaluate the generalization on the test data set. We use 40 and 400 ground state wave functions for the training and test phases, respectively (see [23] for details).
@@
We consider a scenario involving corrupted labels to evaluate the effectiveness of data-based Q-CurL in handling data difficulty during training. With a noise level probability $p$ ($0\leq p\leq 1$), the true label $y_{i}\in\{0,1\}$ of training state |$\ket{\psi_{i}}$ is flipped to the label $1-y_{i}$ with probability $p$, while it remains unchanged with probability $1-p$. Figure 3 illustrates the performance of a trained QCNN on test data across various noise levels. There is a minimal difference at low noise levels, but as noise increases, conventional training fails to generalize effectively. Introducing data-based Q-CurL in training (red lines) reduces test loss and improves test accuracy compared to the conventional method (blue lines). As further presented in  [23], Q-CurL enhances phase separation in the phase diagram, offering more reliable insights into the use of QML for understanding physical systems.